newton ring are formed in reflected light of wavelength 6000A with a liquid between the plane and curved surface .if the diameter of the 6th bright ring is 3.1mm and radius of curvature of curved surface if 100cm calculate the index of the liquid sir answer plz
1
Expert's answer
2018-10-18T11:23:09-0400

For the bright ring system:

μ (d_n^2)/4R=nλ

The refractive index of the liquid

μ=4Rnλ/(d_n^2 )=4(1)(6)(6000∙〖10〗^(-10) )/(0.0031)^2 =1.5